OpenStudy (anonymous):

A sand box has an area of 45ft^2. The length is 4ft longer than the width. Find the dimensions of the sand box. solve by completing the square A. write an equation for the total area. B. find the dimension

OpenStudy (bibby):

what's the area of a rectangle?

OpenStudy (anonymous):

x:width x+4: length area= x*(x+4)=45 x^2+4x-45=0 (x+9)(x-5)=0

OpenStudy (anonymous):

x =-9 or x=5 since there is no area sleek-feathered onetive x=5 width=5 length=5+4=9 area=5*9=45

OpenStudy (anonymous):

thank you